VT004

PI: Edwards
Title: VSOP-Test

Session: Oct 1996 (session 3)
Frequency: 1.3 cm
Mode: VLBA
Obs time (UT): 290 EU 06:00(16/10)-09:00(16/10)
Stations: EbJbNtMhSh
Processor: Soc
EU Mk3 Tapes/St.: 1

Comment


To give feedback on this experiment, click here

Experiment Feedback

Westerbork (Tony Foley at Wed Oct 16 07:27:24 EDT 1996) Minor failures
We could not observe in this mode, but have sent a VLBI tape to Socorro . The tape contains no data.

Noto (Corrado Trigilio at Thu Oct 17 06:24:07 EDT 1996) Success
No problem reported. Weather conditions: cloudy sky, from 30% (at start) up to 80% of coverage. Due to the low elevation of the sources (<30 deg), a strong attenuation of the signal is expected. The tape has been sent to Socorro at the end of the observations (16-Oct-96 - via CAI-post).

Effelsberg (Walter Alef at Fri Oct 18 10:26:49 EDT 1996) Minor failures
Started up in time. Had computer and computer communication problems. Lost some scans partly or completely. Towards the end problems diagnosed and workaround establisched. Last hour or so should be OK. Tape sent off with UPS.

Metsahovi (Kirsi Karlamaa, Ari Mujunen, Kaj Wiik at Sat Oct 19 08:41:02 EDT 1996) Minor failures
(Actually the database does not seem to like single quotes...) The formatter time jumped +1sec at 290-0730 UT; this was detected only later in the log. The schedule left a little bit too little time for slewing in 3--4 passes, maybe this was intentional? There is a suspicious head positioning response of 9629008291423/pass/20,,auto,223.0,,54.7,,-168.3, but the FS did not complain in chekr later. We were running with our old A/D board instead of the new P&G DSM; we would like to test it later in Nov-1996 VSOP test.

Shanghai (Huang Xinyong at Tue Oct 22 03:28:26 EDT 1996) Minor failures
This is the first 22GHz VLBI observation at Shanghai.The Field System has been upgraded to FS9.1.5. The FS System allows generation of SNAP and procedure files. We checked the spectrum at LO frequency, it seems that is OK. Since we just repaired the azimuthal track rail of telescope, the systematic pointing check did not carry out. Only one strong source was observed, because of the bad weather. There is one problem with VLBA recorder. The pressure of vacuum dropped two times during the observation. We try to use software for the generation of the ANTAB calibration, but failed. The tape has been sent to Socorro on 17, Oct., and the log file has been put the ASTOB1 on 18, Oct.

Jodrell (Alastair Gunn at Thu Nov 7 05:52:35 EST 1996) Minor failures
Telescope offsets were not set correctly until 07:45 UT. End of tape detector failed and tape repositioned at 08:26 UT. Weather was wet and windy. Tape IAA00132 shipped to Mitaka on 18/10/96 via DHL.